HETAV PATEL

Grade 11,

Calculate the volume of a 1 mole of an ideal gas at STP

Calculate the volume of a 1 mole of an ideal gas at STP

Grade:11

1 Answers

Mukesh Sharma
10 years ago
here
STP means the gases have pressure 1atm and Temp of 273 Kelvin

Now we're going to use the equation PV= nRT, where P=pressure, V=volume, n= moles, R= constant, T= Temp

First we need to find moles : n= m/MM The Molar Mass of N2= 28.02 moles/gram . The molar mass of O2= 32 moles/gram

Divide the respective masses by the Molar Masses

For N2: n = 73.8/28.02

n=2.63 moles N2

For O2: n= 52.9/32

n= 1.65 moles O2

Use the equation PV= nRT now, in terms of V

V= (nRT)/P

R= constant= 0.0821 (rounded), @STP Temp= 273k and P=1 atm

V of N2 will be:

[ (2.63)(.0821)(273) ] / 1

Volume of N2 = 58.7 Liters

For O2:
[ (1.65)(.0821)(273) ] / 1

Volume of O2 =37.0 Liters

Thanks and regards
askIITians faculty
mukesh sharma